OpenStudy (anonymous):

Celia is staring at the clock waiting for school to end so that she can go to track practice. She notices that the 4-inch-long minute hand is rotating around the clock and marking off time like degrees on a unit circle. Part 1: How many radians does the minute hand move from 1:25 to 1:50? (Hint: Find the number of degrees per minute first.) Part 2: How far does the tip of the minute hand travel during that time? Part 3: How many radians on the unit circle would the minute hand travel from 0° if it were to move 5π inches? Part 4: What is the coordinate point associated with this radian me

OpenStudy (anonymous):

Could somebody please just check the work for me :) Part 1: the number of degrees per minute is 6. The minute hand moves 150 degrees from 1:25-1:50. radians= 150*pi/180 The number of radians is 5pi/6 Part 2: to find the how far the tip of the minute hand travels you would have to find the arc length. s=4*5pi/6= 10pi/3 Part 3: 5pi=4*radians 5pi/4=radians Part 4: 5pi/4*180/pi= 225 degrees (-radical2/2,-radical2/2)